8 u
_ = __
7 11
need to solve for U. These are both fractions

Answer by RAY100(1637)

8/7 = u/11,,,,,,cross multiply
.
8*11 = 7*U,,,,,,divide by 7
.
(8*11)/7 = U
.
88/7 = U
.
12.57 = U
.
.check,,,,8/7 = 12.57 / 11 = 1.142,,,,,ok
.